OpenStudy (anonymous):

if the point (2/3,y) on a unit circle what is y? a) 2/3 b)1/3 c)√5/3 d)4/9 please help!!!!!!

OpenStudy (anonymous):

please help me i can not get it wrong

OpenStudy (kirbykirby):



OpenStudy (kirbykirby):

y is going to be the height of that triangle. ANd you notice you have a right-angled triangle, so you can use the Pythagorean theorem

OpenStudy (kirbykirby):

I drew 1 for the hypothenuse since you have a unit circle, and the radius in a unit circle is 1

OpenStudy (anonymous):

so\[\frac{ 2 }{ 3 } ^{2}\] + \[y ^{2}\] = 1^2

OpenStudy (kirbykirby):

um yes, just to be sure write it as \(\large \left( \frac{2}{3}\right)^2\) All you have to do then is solve for \(y\) :)

OpenStudy (anonymous):

so d

OpenStudy (kirbykirby):

not quite

OpenStudy (kirbykirby):

\[\left( \frac{2}{3}\right)^2+y^2=1^2\\ \frac{4}{9}+y^2 = 1\\ y^2=1-\frac{4}{9}=\frac{5}{9} \]

OpenStudy (kirbykirby):

then square root both sides to find y